# Geometric Freeze-Tag Problem

### v1

@@ 2. Our Results
Finally, we examine a version of FTP in $(\mathbb{R}^{3},l_{2})$, where the asleep robots are located on the boundary of the unit $l_{2}$-ball. Using our approach for $(\mathbb{R}^{2},l_{2})$, we show that FTP in $(\mathbb{R}^{3},l_{2})$ for asleep robots on the boundary can be solved with a makespan of at most $9.26$. Formally we have: {theorem} A robot at the origin can wake up any set of $n$ asleep robots on the boundary of the unit $l_{2}$-ball in $\mathbb{R}^{3}$ with a makespan of at most $9.26$. The proof of this theorem is deferred to Section 5.2.
@@ 3.1.1. Arc-Strategy
$\sum_{i=1}^{k}{d_{i}}=\sum_{i=1}^{k}{\sqrt{b_{i}^{2}+a_{i}^{2}-2a_{i}b_{i}\cos{\beta_{i}}}}$
Since $\beta_{i}=\frac{\pi}{2}-\frac{\alpha_{i}}{2}$, the path length simplifies to
$\sum_{i=1}^{k}{d_{i}}=\sum_{i=1}^{k}{\sqrt{b_{i}^{2}+a_{i}^{2}+2a_{i}b_{i}\sin{\frac{\alpha_{i}}{2}}}}$
Given $b_{i}=2r_{i}\sin{\frac{\alpha_{i}}{2}}$, the total path length becomes:
$\sum_{i=1}^{k}{d_{i}}$ $=$ $\sum_{i=1}^{k}{\sqrt{a_{i}^{2}+4r_{i}^{2}(\sin{\frac{\alpha_{i}}{2}})^{2}+4a_{i}r_{i}(\sin{\frac{\alpha_{i}}{2}})^{2}}}$
@@
$\sum_{i=1}^{k}{\sqrt{a_{i}^{2}+4r_{i}^{2}(\sin{\frac{\alpha_{i}}{2}})^{2}+4a_{i}r_{i}(\sin{\frac{\alpha_{i}}{2}})^{2}}}$
$\leq\sum_{i=1}^{k}{\sqrt{a_{i}^{2}+4r_{i}^{2}(\sin{\frac{\alpha_{i}}{2}})^{2}+4a_{i}r_{i}\sin{\frac{\alpha_{i}}{2}}}}$
$=\sum_{i=1}^{k}{\sqrt{(a_{i}+2r_{i}\sin{\frac{\alpha_{i}}{2}})^{2}}}=\sum_{i=1}^{k}{a_{i}+2r_{i}\sin{\frac{\alpha_{i}}{2}}}$
@@
$\sum_{i=1}^{k}{d_{i}}$$\leq$ $\sum_{i=1}^{k}{a_{i}}+2\sum_{i=2}^{k}{r_{i}\sin{\frac{\alpha_{i}}{2}}}+2r_{1}\sin{\frac{\alpha_{1}}{2}}$
$\sum_{i=1}^{k}{d_{i}}$$\leq$ $1+2\sum_{i=2}^{k}{\sin{\frac{\alpha_{i}}{2}}}+2r_{1}\sin{\frac{\alpha_{1}}{2}}$
@@ 3.1.2. Ring-Strategy
$a_{i}$ $=$ the distance between ($\text{the distance between }(r_{i},\alpha_{i})\text{ and }(r_{i},\alpha_{i+1})$
@@
$b_{i}$ $=$ the distance between ($\text{the distance between }(r_{i},\alpha_{i+1})\text{ and }(r_{i+1},\alpha_{i+1})$
@@
Since we split the half-ring associated with each active point at each step, we have $b_{i}\leq\frac{1-r_{1}}{2^{(i-1)}}$. Thus,
$\sum_{i=1}^{k}{d_{i}}\leq\pi+1-r_{1}+\sum_{i=1}^{k}{\frac{1-r_{1}}{2^{(i-1)}}}$
@@ 4. FTP in $(\mathbb{R}^{3},l_{1})$
The initially awake robot, starting at the center of the cross-polytope, can reach any of the $n$ sleeping robots in $r$ time units. Once the first robot is awakened, both robots can each wake another sleeping robot in $d=2r$ time units. This doubling process continues, with the number of awake robots doubling every $d$ time units. After $6d=12r$ time units, there will be at least $1282\times 2^{6}=$ 128 awake robots. Thus, using this strategy, up to 127 robots can be awakened in $13r$ time units. ∎

### v2

@@ 2. Our Results
Finally, we examine a version of FTP in $(\mathbb{R}^{3},l_{2})$, where the asleep robots are located on the boundary of the unit $l_{2}$-ball. Using our approach for $(\mathbb{R}^{2},l_{2})$, we show that FTP in $(\mathbb{R}^{3},l_{2})$ for asleep robots on the boundary can be solved with a makespan of at most $9.26$. Formally, we have: {theorem} A robot at the origin can wake up any set of $n$ asleep robots on the boundary of the unit $l_{2}$-ball in $\mathbb{R}^{3}$ with a makespan of at most $9.26$. The proof of this theorem is deferred to Section 5.2.
@@ 3.1.1. Arc-Strategy
$\sum_{i=1}^{k}{d_{i}}=d_{1}+\sum_{i=2}^{k}{\sqrt{b_{i}^{2}+a_{i}^{2}-2a_{i}b_{i}\cos{\beta_{i-1}}}}$
Since $\beta_{i}=\frac{\pi}{2}+\frac{\alpha_{i}}{2}$, the path length simplifies to
$\sum_{i=1}^{k}{d_{i}}=d_{1}+\sum_{i=2}^{k}{\sqrt{b_{i}^{2}+a_{i}^{2}+2a_{i}b_{i}\sin{\frac{\alpha_{i-1}}{2}}}}$
Given $b_{i+1}=2r_{i}\sin{\frac{\alpha_{i}}{2}}$, the total path length becomes:
$\sum_{i=1}^{k}{d_{i}}$ $=d_{1}+$ $\sum_{i=2}^{k}{\sqrt{a_{i}^{2}+4r_{i-1}^{2}(\sin{\frac{\alpha_{i-1}}{2}})^{2}+4a_{i}r_{i-1}(\sin{\frac{\alpha_{i-1}}{2}})^{2}}}$
@@
$\sum_{i=2}^{k}{\sqrt{a_{i}^{2}+4r_{i-1}^{2}(\sin{\frac{\alpha_{i-1}}{2}})^{2}+4a_{i}r_{i-1}(\sin{\frac{\alpha_{i-1}}{2}})^{2}}}$
$\leq\sum_{i=2}^{k}{\sqrt{a_{i}^{2}+4r_{i-1}^{2}(\sin{\frac{\alpha_{i-1}}{2}})^{2}+4a_{i}r_{i-1}\sin{\frac{\alpha_{i-1}}{2}}}}$
$=\sum_{i=2}^{k}{\sqrt{(a_{i}+2r_{i-1}\sin{\frac{\alpha_{i-1}}{2}})^{2}}}=\sum_{i=2}^{k}{a_{i}+2r_{i-1}\sin{\frac{\alpha_{i-1}}{2}}}$
@@
$\sum_{i=1}^{k}{d_{i}}$$\leq$ $d_{1}+\sum_{i=2}^{k}{a_{i}}+2\sum_{i=2}^{k}{r_{i-1}\sin{\frac{\alpha_{i-1}}{2}}}$
Note that $\sum_{i=1}^{k}a_{i}=1$ and $a_{1}=d_{1}$ so by replacing $d_{1}$ and considering all $r_{i}\leq 1$ we have:
$\sum_{i=1}^{k}{d_{i}}$$\leq$ $1+2\sum_{2=1}^{k}{\sin{\frac{\alpha_{i}}{2}}}+2r_{1}\sin{\frac{\alpha_{1}}{2}}$
@@ 3.1.2. Ring-Strategy
$a_{i}$ $=$ the distance between ($\text{the distance between }(r_{i-1},\alpha_{i-1})\text{ and }(r_{i-1},\alpha_{i})$
@@
$b_{i}$ $=$ the distance between ($\text{the distance between }(r_{i-1},\alpha_{i})\text{ and }(r_{i},\alpha_{i})$
@@
Since we split the half-ring associated with each active point at each step, we have $b_{i+1}\leq\frac{1-r_{1}}{2^{(i-1)}}$ for $i\geq 1$ and $\sum_{i=1}^{k}a_{i}\leq\pi$. Thus,
$\sum_{i=1}^{k}{d_{i}}\leq\pi+b_{1}+\sum_{i=2}^{k}{\frac{1-r_{1}}{2^{(i-2)}}}$
@@ 4. FTP in $(\mathbb{R}^{3},l_{1})$
The initially awake robot, starting at the center of the cross-polytope, can reach any of the $n$ sleeping robots in $r$ time units. Once the first robot is awakened, both robots can each wake another sleeping robot in $d=2r$ time units. This doubling process continues, with the number of awake robots doubling every $d$ time units. After $6d=12r$ time units, there will be at least $1282\times 2^{6}=$ 128 awake robots. Thus, using this strategy, up to 127 robots can be awakened in $13r$ time units.